OpenStudy (anonymous):

Determine the equation of the line passing through (2, 7) and perpendicular to −8x−7y=−65 .

OpenStudy (amistre64):

perp lines have the same setup you just flip the coeffs for x and y and change negate the operation

OpenStudy (amistre64):

−8x−7y= −65 ; -65 i useless, toss it out -8x -7y = n ; now flip the coeefs -7x -8y = n ; now negate the operation -7x + 8y = n ; now just insert your point to find the right value for "n"